### v1

@@ 3 Discussion
In this work, we propose a deep generative model named DeepBioisostere for designing appropriate bioisosteric replacements to control multiple molecular properties. Our model proceeds a chemical modification in an end-to-end manner, choosing a fragment to remove or insert along with its attachment orientation. We demonstrate diverse scenarios of dual property control, where one property is changed while the other is maintained. DeepBioisostere, in every scenario, successfully modulated their properties in the desired manner by replacing fragments from the given molecules. We further show the adjustability of DeepBioisostere; it dynamically selects which fragment to remove depending on the given property control condition. It is practically advantageous where previous database mining approaches heavily rely on the help of experts for choosing a substructure to be modified. Moreover, we justify the significance of substituents near the replacement site to control comprehensive molecular properties such as drug-likeness, especially for the insertion fragment selection. Our model can design novel bioisosteric replacements that are not included in the training data via flexible selections of fragment pairs. We emphasize that it is particularly important when the removal fragment is uncommon, providing a wide variety of potential chemical modifications. Finally, we applied DeepBioisostere to a practical example of encountering drug-resistant mutation of SARS-CoV-2 main protease. We could obtain potentially improved molecules that strongly interact with the mutant residue while retaining drug-likeness, synthetic accessibility, and size, showcasing the strength of our model.
@@ 4 Methods
4.1 Related works
One of the representative works in database mining approaches is SwissBioisostere.[14, 55] Its bioisostere database is constructed through the matched molecular pair analysis (MMPA) of the ChEMBL database, which conducts the statistical analysis to count the occurrences of each bioisosteric replacement and the resultant changes in biochemical properties, such as log, topological polar surface area (TPSA), and bioactivity in the biochemical assays. SwissBioisostere recommends bioisosteric replacements for a queried molecular fragment based on their frequency in the parent database and expected shifts in the desired properties. The database is easily accessible on the website.
Deep generative modeling has been adopted in molecule optimization, mainly focusing on multiple-property control—the essential goal of the lead optimization process. One major category utilizes a language representation of a molecular structure(SMILES),[56] formulating molecular optimization as a language translation problem.[30, 31, 32] He et al. [30] have utilized natural language as a condition with two machine translation models to optimize absorption, distribution, metabolism, excretion, and toxicity(ADMET). In parallel, Yang et al. [32] have used a constrained transformer together with the state-of-the-art ADMET prediction model, demonstrating successful modifications where core structures of original molecules were retained while achieving desired property enhancements. Although they showed attractive performance in various multi-property control tasks, the overall structure of an original molecule is often not retained in the resultant structures, which might not be preferable in a lead optimization process. Also, since language representation of a molecular structure assumes a particular order of atoms in a molecule, they cannot effectively assess the compatibility between a newly introduced fragment and its surroundings, which is essential for handling complicated molecular properties such as drug-likeness and synthesizability.
An important reference to this work is Chen et al. [35], where molecular optimization is redefined as a graph-editing task via single bond disconnection. Their generative model, Modifier-with-one-fragment(Modof), performs an R-group replacement to achieve a desired property by selecting a bond to be cleaved from an original molecule, detaching a single fragment, and attaching another fragment. Since the Modof model sequentially generates the structural change and applies it to the original structure, it can modify a molecular structure while not changing the residual substructure.
In this work, we attempt to model the overall process of chemical modification through bioisosteric replacements with deep generative modeling, contributing to the autonomous optimization of molecular structure. While database mining approaches, including the SwissBioisostere database, provide bioisosteres for a molecular fragment selected by a medicinal chemist, DeepBioisostere by itself determines which substructure to remove from a target molecule as well as inserts a proper alternative fragment (see Fig. 1). Through its fragment-level modification, DeepBioisostere ensures the overall structure is retained and can replace not only terminal R-groups but also internal substructures of target molecules with their bioisosteres.
@@ 4.2 Finding molecular pairs with bioisosterism: dataset construction
In this section, we present our scheme for constructing a training dataset. It is a pivotal step in our work, where the dataset should be precisely tailored to align with the objective of DeepBioisostere. To obtain a dataset with potential bioisosteric replacements, we begin from the ChEMBL database[27] and go through two steps of training dataset construction: pre-filtration and MMP identification.
@@ 4.2.1 Pre-filtration
We adopted a permissive pre-filtration strategy to maximally retain the data. Simply, we excluded heavy molecules with molecular weights exceeding 800 and molecules consisting of salts. Additionally, molecules with an inhibitory concentration greater than 10 are also excluded so that the remaining molecules can be regarded as possessing sufficient bioactivities. This pre-filtration process provides a refined dataset appropriate to the next stage of MMP identification. From the total of 3,849,629 ChEMBL molecules, we were able to obtain 964,999 molecules by pre-filtration.
4.2.2 Identifying MMPs based on the BRICS
@@
MMP is a pair of molecules in which all parts are identical except for a small fragment in each molecule. From the pre-filtered ChEMBL database, we identified MMPs by fragmenting each molecule and then selecting molecule pairs that share common fragments. It is crucial to adopt a proper fragmentation rule in this process since it determines the fragments corresponding to the difference between each MMP, which are to be considered potential bioisosteres. To consider more synthetically accessible modifications than the conventional strategy of simply cleaving rotatable bonds in the fragmentation process, we employed BRICS[36] rules for molecular fragmentation, implemented in the RDKit[57] package.
@@
Using the BRICS rules, the ChEMBL molecules are represented as fragment-level graphs (see Fig. 6(A)). Each node is irreducible fragment, a subgraph of a molecule that cannot be further divided via fragmentation, where edges between nodes correspond to the BRICS bonds. In this regard, a removal fragment is defined as a single subgraph of a fragment-level molecular graph. With these definitions, we enumerate multiple fragment combinations that involve a single removal fragment and remaining fragments for each BRICS-fragmented molecule. The fragment combinations are further filtered with two criteria about their removal fragments: (1) a maximum of 12 heavy atoms in a removal fragment and (2) the removal fragment should have fewer heavy atoms than the remaining fragments.
@@
The aforementioned process results in an enormous number of fragment combinations derived from diverse molecules. Then, the obtained fragment combinations are subsequently compared, where any molecular pairs with the same residual fragments are classified as MMPs along with their removal fragments. We note that the majority of the resultant MMPs contain more than two matched fragment combinations. For these cases, we exclusively selected the smallest removal fragments as bioisosteres (see Fig. 6(B)) This selection of the smallest removal fragment distinguishes our training dataset from previous database mining approaches that enumerate every bioisosteric pair, preventing the accumulation of redundant data in the training data. The final training dataset consists of a total of 24,163,198 MMPs specified about their corresponding bioisosteres.
@@
Based on the final dataset, we collect the insertion fragments to build a fragment library. The fragment library is necessary for both the training and inference phases of DeepBioisostere as a source of insertion fragment candidates, which our model can utilize to optimize a given molecule. For more details on the use of the fragment library of DeepBioisostere, refer to the Sections 4.4 and 4.5. Note that gathering all removal fragments will result in the same fragment library since each identified MMP results in two training data entries with reverse relationships in their removal and insertion fragments. The resulting fragment library is composed of 206,008 diverse fragments. Furthermore, we randomly split the fragment library at a ratio of 8:1:1, leading to 164,813, 20,604, and 20,591 fragments for training, validation, and test datasets, respectively.
@@ 4.3 DeepBioisostere: Deep generative modeling of bioisosterism
where is a removal fragment, is an insertion fragment, and is an attachment orientation. Note that for a modification of to , there might be more than one set of the three modification components, but we uniquely defined and to be only the smallest removal and insertion fragments in the training data construction procedure, respectively. In this way, the ground truth for is defined to be zero if and are not the smallest ones to modify to another molecular structure, . Since the smallest structural change can always be identified from any MMP, this choice does not impose any restriction on the probability distribution of . The details about how we model each conditional probability of the Eq. 1 are discussed in the method section.
@@ 4.3.1 Fragment-level molecular encoding
The overall framework of DeepBioisostere is depicted in Fig. 7. First, we encode an original molecule to be modified into fragment-level embedding vectors. We adopt a hierarchical architecture to facilitate probability modelings on fragment-level information afterward. An original molecule is represented in two ways: an atom-level graph, , and a fragment-level graph, . Largely inspired by Chen et al. [35], we employ two message-passing neural networks in different levels named an atom message passing network(AMPN) and a fragment message passing network(FMPN), respectively. An original molecule is first encoded into atom-level embedding vectors by AMPN. As illustrated in Fig. 7(A), AMPN runs within the whole graph and subgraphs of a molecule, which adopts respective atom-level embedding vector update mechanisms and uses differently initialized weights. An initial atom feature vector, , is concatenated with the condition vector, C, to yield an input atomic embedding vector by . Details about constructing the initial atom feature vectors are summarized in the supplementary information. We summarized key notations for model architecture in Table 2. Embedding vectors for both atoms and bonds are propagated through the AMPN layer as follows:
@@
where is neighboring atoms of , GRU is the gated recurrent unit introduced in Chung et al. [58] The whole graph AMPN operates on the entire molecular graph , updating atom-level embeddings based on the collective information of all atoms and bonds in the molecule. In contrast, the subgraph AMPN works locally within each irreducible fragment, focusing its update on atoms only within the specific fragment, denoted as . This subgraph-based update strategy allows the model to learn the comprehensive features of each irreducible fragment for further hierarchical message passing.
Notation | Description |
Atom-level molecular graph | |
, | Set of atoms and bonds in , respectively |
An atom in | |
A bond connecting atoms and in | |
An embedding vector of in layer | |
An embedding vector of | |
Fragment-level molecular graph | |
, | Set of irreducible fragments and edges in , respectively |
A subgraph, fragmented by the BRICS rules | |
An irreducible fragment | |
A removal fragment | |
An insertion fragment | |
An edge between irreducible fragments and in , corresponding to a BRICS bond in | |
A set of atoms included in a subgraph | |
A set of atoms composing | |
A set of irreducible fragments included in a subgraph | |
An embedding vector of in layer | |
An embedding vector of | |
An aggregated vector for | |
An attachment orientation, a set of edges enumerated based on the BRICS rules | |
Element-wise concatenation | |
Sigmoid activation function | |
@@ 4.3.2 Modeling sequential modification steps
Similarly, is an insertion fragment selection module parameterized with . Thus, the insertion fragment depends on the pre-selected removal fragment, its surroundings, and the given property control condition, which are comprehensively included in .
@@ 4.4 Training DeepBioisostere
where , , is a allowed but negative fragment from the original molecule, and is the total number of allowed but negative fragments from the original molecule.
The selection of insertion fragments is more complicated since our model should be able to retrieve the ground truth one among a vast number of potential ones from the fragment library. To avoid the inefficiency in evaluating tens of thousands of fragments for every data point, we employed the negative sampling strategy inspired by Seo et al. [59] This strategy samples only a few fragments (20 fragments in this work) with the probability weakly proportional to their occurrence as negative samples. With this negative sampling strategy, we formulated as follows:
where , , is the number of negative samples for each positive data. By including negative samples in Eq. 13, we counteracted the bias towards always preferring more frequently occurring fragments over less frequent ones.
@@
where and . and are the number of potential edges within the positive and negative orientations, respectively.
For training and validation, we split the whole dataset constructed in Section 4.2 into training, validation, and test sets with a ratio of 8:1:1. The dataset was split according to their insertion fragments so that the three sub-sets share no insertion fragment. Thus, each dataset contains distinct bioisosteric replacements. For the multi-property control, we used a total of four properties that can be readily computed based on the structure of a molecule: molecular weight, log, quantitative estimate of drug-likeness(QED),[39] and SAscore.[40] The training was finished in no more than 17 hours with a single NVIDIA RTX A4000 GPU for each property control condition. Details about the training process can be found in the supplementary information.
@@ 4.5 Molecular Optimization Process with DeepBioisostere
For each selected removal fragment, DeepBioisostere explores the fragment library to find suitable insertion fragments for property control. Since the chemical modification of a certain part should retain its surroundings, molecular fragments with different numbers of fragmented sites are not considered. In addition, among the corresponding fragments, only those that satisfy the BRICS rules at the respective removal site of the original molecule are considered. All the passed insertion fragments are then evaluated by the DeepBioisostere model, and top insertion fragments are greedily selected. We note that, during the insertion fragment selection, either their occurrence data from the MMP database we created or from public chemical libraries like ChEMBL is not utilized because the differences in their occurrences are implicitly captured during the training process.
Lastly, the selected insertion fragments are examined in their attachment orientation. The combinations between fragmented sites of the insertion fragment and the surroundings are enumerated and again filtered with the BRICS rules. The passed combinations are then prioritized by the DeepBioisostere model and the insertion fragments are attached following the selected attachment orientations. More details about the generation process can be found in the supplementary information.
@@
The code for processing data from the ChEMBL database is available at Github: https://github.com/Hwoo-Kim/DeepBioisostere.git
@@
The code for training DeepBioisostere, sampling molecules, and evaluating the generated structures is available at Github: https://github.com/Hwoo-Kim/DeepBioisostere.git

### v2

@@ 3 Discussion
In this work, we propose a deep generative model named DeepBioisostere for autonomous bioisosteric replacement with fine control over multiple molecular properties. Our model conducts chemical modification in an end-to-end manner, from selecting fragments to remove and insert, along with their attachment orientation. We demonstrate various scenarios of dual property control, where one property is altered while the other remains unchanged. DeepBioisostere successfully modulated the properties of the given molecules in all scenarios by replacing the fragments in the desired manner. We further demonstrate the adjustability of DeepBioisostere, which dynamically selects the fragment to remove based on the given property control condition. This is practically advantageous, where previous database-mining approaches heavily rely on the help of experts for choosing a substructure to be modified. Moreover, we justify the significance of substituents near the replacement site in controlling comprehensive molecular properties, especially for the selection of insertion fragments, which could not be addressed by conventional in silico approaches. Our model enables more diverse replacements appropriate for the given property condition, extending from the existing moiety pairs within the database. We emphasize that this is particularly important when the removal fragment is uncommon, as it provides a wide variety of potential chemical modifications. Finally, we applied DeepBioisostere to a practical scenario that mimics the hit-to-lead process, which involves optimizing the drug-likeness and synthetic accessibility of computationally generated hit compounds while retaining their binding affinities. These results demonstrated that DeepBioisostere successfully tuned multiple properties without disrupting binding affinity for all 3D molecular generative models used, showcasing the promising role of our approach in autonomous drug design.
While DeepBioisostere already demonstrates the ability to control multiple molecular properties through a deep generative framework, its design also opens promising directions for further enhancement. In particular, extending the model with multi-task and transfer learning could enable it to handle conditions with limited or noisy labels more effectively. Pretraining on widely available or easy-to-label molecular properties, such as QED, logP, and SA score, and transferring the learned representations to tasks involving complex endpoints, like ADMET properties, would enable the framework to generalize across a broader chemical and biological space. Such advancements could further strengthen the model’s applicability to diverse drug discovery pipelines, facilitating more accurate and data-efficient molecular design in increasingly realistic and challenging scenarios.
@@ 4.1 Finding Molecular Pairs with Bioisosterism: Dataset construction
In this section, we present our scheme for constructing a training dataset. It is a pivotal step in our work, where the dataset should be precisely tailored to align with the objective of DeepBioisostere. To obtain a dataset with potential bioisosteric replacements, we begin from the ChEMBL database[ChEMBL] and go through two steps of training dataset construction: pre-filtration and MMP identification.
@@ 4.1.1 Pre-filtration
We adopted a permissive pre-filtration strategy to maximally retain the data. Simply, we excluded heavy molecules with molecular weights exceeding 800 and molecules consisting of salts. Additionally, molecules with an inhibitory concentration greater than 10 are also excluded so that the remaining molecules can be regarded as possessing sufficient bioactivities. This pre-filtration process provides a refined dataset appropriate to the next stage of MMP identification. From the total of 4,182,178 ChEMBL molecules, we were able to obtain 1,054,335 molecules by pre-filtration.
4.1.2 Identifying Bioisosteric Replacements based on the BRICS MMPs
@@
MMP is a pair of molecules in which all parts are identical except for a small fragment in each molecule. From the pre-filtered ChEMBL database, we identified MMPs by fragmenting each molecule and then selecting molecule pairs that share common fragments. It is crucial to adopt a proper fragmentation rule in this process since it determines the fragments corresponding to the difference between each MMP, which are to be considered potential bioisosteres. To consider more synthetically accessible modifications than the conventional strategy of simply cleaving rotatable bonds in the fragmentation process, we employed BRICS[BRICS] rules for molecular fragmentation, implemented in the RDKit[rdkit] package.
@@
Using the BRICS rules, the ChEMBL molecules are represented as fragment-level graphs (see Fig. 6(a)). Each node is irreducible fragment, a subgraph of a molecule that cannot be further divided via fragmentation, where edges between nodes correspond to the BRICS bonds. In this regard, a removal fragment is defined as a single subgraph of a fragment-level molecular graph. With these definitions, we enumerate multiple fragment combinations that involve a single removal fragment and remaining fragments for each BRICS-fragmented molecule. The fragment combinations are further filtered with two criteria about their removal fragments: (1) a maximum of 12 heavy atoms in a removal fragment, and (2) the removal fragment should have fewer heavy atoms than the remaining fragments.
@@
The aforementioned process results in an enormous number of fragment combinations derived from diverse molecules. Then, the obtained fragment combinations are subsequently compared, where any molecular pairs with the same residual fragments are classified as MMPs along with their removal fragments. We note that the majority of the resultant MMPs contain more than two matched fragment combinations. For these cases, we exclusively selected the smallest removal fragments as bioisosteres (see Fig. 6(b)) This selection of the smallest removal fragment distinguishes our training dataset from previous database-mining approaches that enumerate every bioisosteric pair, preventing the accumulation of redundant data in the training data. To ensure that only MMPs preserving bioactivity were considered, pairs with were filtered out. The final training dataset consists of a total of 2,874,580 MMPs specified about their corresponding bioisosteres.
@@
Based on the final dataset, we collect the insertion fragments to build a fragment library. The fragment library is necessary for both the training and inference phases of DeepBioisostere as a source of insertion fragment candidates, which our model can utilize to optimize a given molecule. For more details on the use of the fragment library of DeepBioisostere, refer to the Sections 4.3 and 4.4. Note that gathering all removal fragments will result in the same fragment library since each identified MMP results in two training data entries with reverse relationships in their removal and insertion fragments. The resulting fragment library is composed of 140,096 diverse fragments. Furthermore, we randomly split the fragment library at a ratio of 8:1:1, leading to 112,076, 14,013, and 14,007 fragments for training, validation, and test datasets, respectively.
@@ 4.2 DeepBioisostere: Deep Generative Modeling of Bioisosterism
where is a removal fragment, is an insertion fragment, and is an attachment orientation. Note that for a modification of to , there might be more than one set of the three modification components, but we uniquely defined and to be only the smallest removal and insertion fragments in the training data construction procedure, respectively. In this way, the ground truth for is defined to be zero if and are not the smallest ones to modify to another molecular structure, . Since the smallest structural change can always be identified from any MMP, this choice does not impose any restriction on the probability distribution of . The details about how we model each conditional probability of the Eq. 1 are discussed in Section 4.2.2.
@@ 4.2.1 Fragment-level Molecular Encoding
The overall framework of DeepBioisostere is depicted in Fig. 7. First, we encode an original molecule to be modified into fragment-level embedding vectors. We adopt a hierarchical architecture to facilitate probability modeling at the fragment level. An original molecule is represented in two ways: an atom-level graph and a fragment-level graph .
We adopted the core architectural design of Modof proposed by Modof, which demonstrates strong molecular property conditioning performance through a hierarchical framework. In its design, as illustrated in Fig. 7, atom-wise embeddings are first generated using an atom-level graph message passing network, then aggregated into fragment-level hidden vectors, which are subsequently processed by a fragment-level graph message passing network. In DeepBioisostere, these two message passing networks are adapted for bioisosteric replacement and are referred to as the atom message passing network (AMPN) and fragment message passing network (FMPN), respectively.
To tailor the model for bioisosteric replacement, we design the AMPN with two submodules having distinct parameters: a whole-graph AMPN operating on all molecular nodes and a subgraph AMPN operating only within individual fragments. The subgraph AMPN is employed as a shared fragment embedder for both the original molecule and the fragment library, providing a consistent and comparable representation of molecular fragments. This weight-sharing strategy ensures a comprehensive understanding of fragments and, in turn, allows the overall generative model to be formulated as a product of fragment-level conditional probabilities (Eq. 1). In contrast to Modof, which decodes new fragments independently from separate latent vectors, DeepBioisostere evaluates candidate replacements by directly comparing the joint embedding of the removal fragment and each candidate insertion fragment from its library in this shared space. Furthermore, when combined with our same-bioassay molecular pair dataset, this molecular encoding strategy enables the model to capture similarity in bioactivity as well as differences in other molecular properties. We quantitatively analyze the contribution of this architectural design to multi-property optimization in Supplementary Information Section 5.1.
An initial atom feature vector, , is concatenated with the condition vector, C, to yield an input atomic embedding vector by . Details about constructing the initial atom feature vectors are summarized in the Supplementary Information. We summarized key notations for model architecture in Table 3. Embedding vectors for both atoms and bonds are propagated through the AMPN layer as follows:
@@
where is neighboring atoms of , GRU is the gated recurrent unit introduced in GRU The whole graph AMPN operates on the entire molecular graph , updating atom-level embeddings based on the collective information of all atoms and bonds in the molecule. In contrast, the subgraph AMPN works locally within each irreducible fragment, focusing its update on atoms only within the specific fragment, denoted as . This subgraph-based update strategy allows the model to learn the comprehensive features of each irreducible fragment for further hierarchical message passing.
| Notation | Description | ||
| Atom-level molecular graph | |||
| , | Set of atoms and bonds in , respectively | ||
| An atom in | |||
| A bond connecting atoms and in | |||
| An embedding vector of in layer | |||
| An embedding vector of | |||
| Fragment-level molecular graph | |||
| A subgraph, fragmented by the BRICS rules | |||
| An irreducible fragment | |||
| A removal fragment | |||
| An insertion fragment | |||
| A set of atoms included in a subgraph | |||
| A set of atoms composing | |||
| A set of irreducible fragments included in a subgraph | |||
| An embedding vector of in layer | |||
| An embedding vector of | |||
| An aggregated vector for | |||
| Element-wise concatenation | |||
| Sigmoid activation function | |||
@@ 4.2.2 Modeling Sequential Modification Steps
Similarly, is an insertion fragment selection module parameterized with . Thus, the insertion fragment depends on the pre-selected removal fragment, its surroundings, and the given property control condition, which are comprehensively included in . This contrastive formulation enables the model to explicitly learn bioisosteric relationships between fragments.
@@ 4.3 Training DeepBioisostere
where , , is an allowed but negative fragment from the original molecule, and is the total number of allowed but negative fragments from the original molecule.
The selection of insertion fragments is more complicated since our model should be able to retrieve the ground truth one among a vast number of potential ones from the fragment library. To avoid the inefficiency in evaluating tens of thousands of fragments for every data point, we employed the negative sampling strategy inspired by SeoBBAR This strategy samples only a few fragments (20 fragments in this work) with the probability weakly proportional to their occurrence as negative samples. With this negative sampling strategy, we formulated as follows:
where , , is the number of negative samples for each positive data. By including negative samples in Eq. 13, we counteracted the bias towards always preferring more frequently occurring fragments over less frequent ones. One potential concern is that, for a given molecular structure and its removal fragment pair, if there are two distinct ground-truth insertion fragments, one could mistakenly be sampled as a negative fragment during training. However, since the final training objective is defined as the expectation over the entire dataset, the model will ultimately assign equal probabilities to both fragments.
@@
where and . and are the numbers of potential edges within the positive and negative orientations, respectively.
For training and validation, we split the whole dataset constructed in Section 4.1 into training, validation, and test sets with a ratio of 8:1:1. The dataset was split according to their insertion fragments so that the three sub-sets share no insertion fragment. Thus, each dataset contains distinct bioisosteric replacements. For the multi-property control, we used a total of four properties that can be readily computed based on the structure of a molecule: molecular weight, log, QED,[qed] and SAscore.[sascore] The training was finished in no more than 17 hours with a single NVIDIA RTX A4000 GPU for each property control condition. Details about the training process can be found in the Supplementary Information.
@@ 4.4 Molecular Optimization Process with DeepBioisostere
For each selected removal fragment, DeepBioisostere explores the fragment library to find suitable insertion fragments for property control. Since the chemical modification of a certain part should retain its surroundings, molecular fragments with different numbers of fragmented sites are not considered. In addition, among the corresponding fragments, only those that satisfy the BRICS rules at the respective removal site of the original molecule are considered. All the passed insertion fragments are then evaluated by the DeepBioisostere model, and the top insertion fragments are greedily selected. We note that, during the insertion fragment selection, either the occurrence data from the MMP database we created or from public chemical libraries, such as ChEMBL, is not utilized because the differences in occurrence are implicitly captured during the training process.
Lastly, the selected insertion fragments are examined in their attachment orientation. The combinations between the fragmented sites of the insertion fragment and its surroundings are enumerated and then filtered using the BRICS rules. The passed combinations are then prioritized by the DeepBioisostere model, and the insertion fragments are attached following the selected attachment orientations. For more details about the generation process, refer to the Supplementary Information.
@@
The code for processing data from the ChEMBL database is available on GitHub: https://github.com/Hwoo-Kim/DeepBioisostere.git
@@
The code for training DeepBioisostere, sampling molecules, and evaluating the generated structures is available on GitHub: https://github.com/Hwoo-Kim/DeepBioisostere.git